# Predicting Trust In Autonomous Vehicles: Modeling Young Adult Psychosocial Traits, Risk-Benefit Attitudes, And Driving Factors With Machine Learning

### v1

@@ 5.2. Limitations and Future Work
This study was not without limitations. First, our study leveraged survey methods, which – though standard practice – may be subject to bias as it is based on self-reported data. To mitigate this concern, we employed a large sample size, used validated questionnaires, and replicated variable selection from prior work whenever possible. Self-report concern may be particularly strong with measurements of trust, as (though quite common) self-reported trust may not be as effective as measuring trust via behavioral reliance (Hoff and Bashir, 2015). To mitigate this risk, we used a composite trust measure to capture many aspects of a trust judgement. Future work may seek to use behavioral trust measures. We focused on young adults, a majority of whom were enrolled at an undergraduate institution in the United States. As such, results presented in this study may not widely generalize. Future work should replicate on other populations if results in other specific settings are sought. Finally, testing the design implications through actual implementation of the guidelines outlined in Section 5.1 will be necessary to assure their validity.
@@ Acknowledgements.
We would like to thank Rohan Bhide, Saumitra Sapre, Chloe Lee, and Janzen Molina for their help disseminating the survey, and David Danks for his advice on modeling methods.

### v2

@@ 5.2. Limitations and Future Work
This study was not without limitations. First, our study leveraged survey methods, which – though standard practice – may be subject to bias as it is based on self-reported data. To mitigate this concern, we employed a large sample size, used validated questionnaires, and replicated variable selection from prior work whenever possible. Self-report concern may be particularly strong with measurements of trust, as (though quite common) self-reported trust may not be as effective as measuring trust via behavioral reliance (Hoff and Bashir, 2015). To mitigate this risk, we used a composite trust measure to capture many aspects of a trust judgement. Variability in participants’ familiarity with AV technology may introduce latent confounds, as those with more exposure might respond differently to questions about AV trust, risks, or benefits. However, we believe this limitation is minimized by focusing on a clear, every day commuting context, which provided a shared baseline for interpretation. This range in familiarity also reflects the diversity of real-world users, enhancing the generalizability of our findings. Future work may segment participants by familiarity levels to help control for baseline knowledge differences. Likewise, we recognize that institutional trust measures like “trust in tech companies” may be interpreted as broad constructs, encompassing varying levels of trust that could be influenced by specific companies or historical practices. While these widely used variables capture overarching attitudes toward technology and institutions, which are valid for the predictive task in this study as they reflect general perceptions relevant to user trust in AVs, future work could benefit from a more nuanced approach to provide clearer insights. For example, exploring sub-features within each institutional construct could help differentiate the effects of brand or institutional identity, transparency practices, or innovation history. Future work may seek to use behavioral trust measures. We focused on young adults, a majority of whom were enrolled at an undergraduate institution in the United States. As such, results presented in this study may not widely generalize. Future work should replicate on other populations if results in other specific settings are sought. Finally, testing the design implications through actual implementation of the guidelines outlined in Section 5.1 will be necessary to assure their validity.
@@ Acknowledgements.
We would like to thank Rohan Bhide, Saumitra Sapre, Chloe Lee, and Janzen Molina for their help disseminating the survey. A special thanks to David Danks for his advice on modeling methods and to the Discovery Way Foundation for their generous usage of lab space.